consider the plane x+2y-3z=2 and 15x-9y-z=2 are they parallel or perpendicular ?or neither parallel or perpendicular?
Answer by ikleyn(52784)

.
The normal vector to the plane x + 2y -3z = 2 is the vector u = (1,2,-3).


The normal vector to the plane 15x - 9y -z = 2 is the vector v = (15,-9,-1).



The vectors u and v are perpendicular, since their scalar product is 0: 1*15 + 2*(-9) + (-3)*(-1) = 0.

Solved.